A company produces and sells a consumer product and is able to control the demand for the product by varying the selling price. The approximate relationship between price and demand is

p = $38 +2,700/D −5,000/D2, for D > 1,

where p is the price per unit in dollars and D is the demand per month.

The company is seeking to maximize its profit. The fixed cost is $1,000 per month and the variable cost (Cv) is $40 per unit.

a. What is the number of units that should be produced and sold each month to maximize profit?

b. Show that your answer to Part (a) maximizes profit


1
Expert's answer
2020-11-23T10:19:02-0500

a. Profit = Revenue – Cost

Cost = fixed cost + (variable cost)D = 1000 + 40D

Revenue = PD = (38 + "\frac{2700}{D}""\frac{5000}{D^2}" )D

Profit = 38D + 2700 - "\frac{500}{D}" – 1000 – 40D

P = 1700 – 2D - "\frac{5000}{D}"

dP/dD = -2 + "\frac{5000}{D^2}"

For maximize profit

dP/dD = 0

D2 = 2500

D = 50

Demand to maximize to profit is 50 units.

b. At D = 49

Profit = 1700 – 2D - "\frac{5000}{49}" = 1499.95

At D = 51

Profit = 1700 – 100 - "\frac{5000}{51}" = 1499.96

At D = 50

Profit = 1700 – 100 - "\frac{5000}{50}" = 1500

So, D = 50 maximize profit.